3. Rolling a Die: If a die is rolled one time, find these probabilities. 

a. Getting a 2 

b. Getting a number greater than 6 

c. Getting an odd number 

d. Getting an odd or even number


1
Expert's answer
2022-06-06T15:58:58-0400

When we roll a die, we can obtain six possible outcomes: 1, 2, 3, 4, 5, 6.

The probability of getting one of these outcomes equals "\frac{1}{6}."


(a) "P(X=2)=\frac{1}{6}."


(b) "P(X>6)=0," this is impossible event.


(c) "P(X \: is \: odd)=P(X\in\{1, 3, 5\})=\frac{3}{6}=\frac{1}{2}."


(d) "P(X \: is \: odd \: or \: even)=P(X\in\{1, 3, 5\})+P(X\in\{2, 4, 6\})=\frac{3}{6}+\frac{3}{6}=1."


Answer: (a) "\frac{1}{6}" (b) 0 (c) "\frac{1}{2}" (d) 1